OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution. 3x + 4y = -2 and 2x - 3y = 10

Parth (parthkohli):

1) Multiply equation 1 by 2. 2) Multiply equation 2 by 3. 3) Subtract both the equations. 4) Solve for y. 5) After solved, substitute the value of y in any of the equations and solve for x.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 6x + 8y = -4 \text{ and } 6x - 9y = 30 }\) Subtract both \(\Large \color{MidnightBlue}{\Rightarrow 17y = -34 }\) \(\Large \color{MidnightBlue}{\Rightarrow y = -2 }\) Substitute: \(\Large \color{MidnightBlue}{\Rightarrow 6x - 16 = -4 }\) \(\Large \color{MidnightBlue}{\Rightarrow 6x = 12 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 6 }\)

OpenStudy (anonymous):

3x+4y=-2 2x-3y=10 Multiply first equation by 3 and second equation by4 3(3x+4y)= 3*-2 4(2x-3y)=4*10 9x+12y=-6 8x-12y=40 Add equation 1 and 2 9x+8x+12y-12y=40-6 17x=34 x=34/17 x=2 Put x=2 in second equation 2x-3y=10 4-3y=10 -3y=10-4 -3y=6 y=6/-3 y=-2 therefore x=2 and y=-2

Parth (parthkohli):

Oops, I meant x = 2

OpenStudy (anonymous):

Thanks (: